Grade 12Organic Chemistry

The vapour pressure of pure liquid A is 10 torr and at the same temperature when 1g of B solid is dissolved in 20g of A,its vapour pressure is reduced to 9 torr if the molecular mass of A is 200amu then the molecular mass of B is

Krisha
8 Years agoGrade 12

3 Answers

Sanam Bedi
8 Years ago
The Actual formula is 
(po-ps)/po=n/(n+N)
 
But for our convinience we will neglect ‘n’ here
(po-ps)/po=n/(N)
 
(10-9)10=(1/x)/(1/10)
(1/10)=(10/x)
(1/100)=(1/x)
x=100amu
 
Therefore, x=100amu
Pooja
7 Years ago
If we will not neglect (n) then answer will be 90 and we must not neglect (n) because it is not mention in question that the amount of A is  more than B 
Kushagra Madhukar
6 Years ago
Dear student,
Please find the solution to your problem below.
 
The Actual formula is 
(po-ps)/po=n/(n+N)
 
(10 – 9)/10 = (1/x) / [(1/x) + (1/10)]
1/10 = 1/(1 + x/10)
1 + x/10 = 10
x/10 = 9
x = 90amu
Therefore, x = 90amu
 
Hope it helps.
Thanks and regards,
Kushagra